Write the quadratic function in vertex form.

y = x2 + 16x + 74

help and show work
1
Expert's answer
2014-06-02T05:33:45-0400
The vertex form of a quadratic function is given by
f (x) = a(x - h)2 + k, where (h, k) is the vertex of the parabola. Therefore we get
y=x2+16x+74=x2+16x+64+10
Using formula (a+b)2=a2+2ab+b2 we obtain
y=(x+8)2+10.